### v1

# Multilayer Network of Cardiovascular Diseases and Depression via Multipartite Projection

Abstract
There is a significant comorbidity between cardiovascular diseases (CVD) and depression that is highly predictive of poor clinical outcome. Yet, its underlying biological pathways remain challenging to decipher, presumably due to its non-linear associations across multiple mechanisms. Mutual information provides a framework to analyze such intricacies. In this study, we proposed a multipartite projection method based on mutual information correlations to construct multilayer disease networks. We applied the method to a cross-sectional dataset from a wave of the Young Finns Study. This dataset assesses CVD and depression, along with related risk factors and two omics of biomarkers: metabolites and lipids. Instead of directly correlating CVD-related phenotypes and depressive symptoms, we extended the notion of bipartite networks to create a multipartite network that connects these phenotype and symptom variables to intermediate biological variables. Projecting from these intermediate variables results in a weighted multilayer network, where each link between CVD and depression variables is marked by its ‘layer’ (i.e., metabolome or lipidome). Using this projection method, we identified potential mediating biomarkers that connect CVD to depression. These included metabolites such as creatinine, valine, phospholipids in very large high-density lipoproteins HDL, triglycerides in small low-density lipoproteins (LDL), and apolipoprotein B; and lipids such as specific sphingomyelins (SM), phosphatidylcholines (PC), triacylglycerols (TAG) and diacylglycerols (DAG). These biomarkers thus may play significant roles in the biological pathways of CVD-depression comorbidity. Additionally, the projected network highlights sex and BMI as the most important risk factors, or confounders, associated with the comorbidity. Our method can generalize to any number of omics layers and disease phenotypes, offering a truly system-level overview of biological pathways contributing to comorbidity.
keywords:
Mutual information, Multilayer disease networks, Multipartite projection, CVD-depression comorbidity, Risk factors, Metabolites, Lipids
Introduction
Cardiovascular diseases (CVD) and depression are pervasive health challenges. The significant co-occurrence of these conditions indicates a profound comorbidity [1, 2, 3]. Indeed, individuals with depression face an elevated risk of developing CVD compared to the general population [4]. Conversely, individuals with CVD are more prone to developing depression, and their comorbidity is linked to a 3-fold higher mortality rate [5, 6]. In either direction, both behavioral and biological factors play a role [5, 7, 8]. An extensive literature documents possible biological mechanisms underlying the bidirectional relationship between CVD and depression, such as increased inflammation, endocrine disturbances, and imbalances in neurotransmitter levels, sympathetic and parasympathetic nervous systems [9, 10, 11, 12]. Either phenotype alone already significantly diminishes the quality of life, imposes a considerable burden on the healthcare system, and remains a challenge to treat —their combination presents an even bigger challenge. Success in solving this major public health concern, e.g., developing effective preventive and treatment strategies, hinges on accurately identifying potential pathways that connect CVD and depression. However, the precise nature of these interactions has not yet clearly been established. Continued effort to understanding their biological associations, by identifying molecular biomarkers and uncovering underlying risk factors, remains important. In this article, we contribute to this effort by introducing a novel disease network construction method, investigating the relationships between CVD-related phenotypes and depressive symptoms, and their biological mediators.
The concept of ’human disease networks’ has been introduced for this purpose [13], providing a powerful framework to study human diseases and underlying biological mechanisms using a network perspective. As an illustrative example, Goh et al. introduced a projection method to reconstruct disease networks based on a bipartite graph [13]. In their projected networks, the association between two diseases was established if there was a gene that was implicated in both. The strength of gene-disease associations was quantified by (linear) Pearson correlation, and disease-disease associations (i.e., explaining their comorbidity) were defined as the number of genes that were significantly associated with both diseases. However, there were several limitations of this methodology that might lead to bias in estimating disease associations. Firstly, it used only genes to connect diseases, whereas in reality there are multiple omics (e.g., genes, proteins, lipids, etc.) connecting diseases. Secondly, simply counting the number of shared intermediate biomarkers was a coarse measure that treated each biomarker equally, no matter their strength of (Pearson) correlation. Thirdly, the use of Pearson correlation means that potentially nonlinear correlations were missed [14]. Lastly, these methods did not allow for the incorporation of higher-order correlations into the projection.
Some studies have expanded on the methodology to address the bias. Park, J. et al. implemented cellular networks and incorporated genes and proteins into their analyses  [15]. To tackle the second limitation, Zhou, X. et al. applied the cosine similarity of respective symptom vectors of given diseases to quantify disease associations, and defined the order of interactions based on the length of path between proteins [16]. Grosdidier, S. et al. proposed a concept of molecular comorbidity index (MCI) based on proteins common between two diseases to estimate the strength of the association [17]. Although extending to multiple omics or defining a more complicated disease association, methodological shortcomings remain. Specifically, none of the current methods provide a combined solution that incorporates multiple omics datasets into a single network analysis, integrates a more refined counting of mediating biomarkers, and utilizes an expressive non-linear correlation.
To fill the above gap, we propose a multipartite projection method that (i) can combine any number of omics datasets and diseases; (ii) is based on the non-parametric Shannon mutual information (MI) correlation; and (iii) counts shared biomarkers and risk factors in a weighted manner. MI is a non-parametric metric that captures both linear and nonlinear mutual interdependence between two variables [18]. Counting biomarkers in a weighted way provides a finer-grained measure to quantify the links between diseases. For this purpose, we propose three definitions for the projected disease links. These definitions not only count the number of their shared neighboring (intermediate) biomarkers, but also consider the strength of their nonlinear correlations with these biomarkers, the importance of these correlation links, and their own neighboring biomarkers that have the shortest path. Compared to direct correlation, the links implied by the network projection provides deeper insights into potential common etiologies.
We applied the projection method to a dataset from the Young Finns Study (YFS) and selected a projection definition that performs better. The dataset encompasses three sets of variables: (1) CVD and depression phenotypes (see Table 1 and  2) and related risk factors (e.g., covariates, Table 4), (2) metabolites, and (3) lipids. Our method and its resulting weighted network allow us to: (i) in parallel identify important biomarkers, representing multiple biological layers possibly related to the development of CVD-depression comorbidity; (ii) within this network, integrate and account for important risk factors; (iii) identify biomarkers related to specific risk-disease and disease-disease associations. The projection method detects sex and body mass index (BMI) as two important risk factors related to both CVD and depression, and identifies significant mediating biomarkers underlying their comorbidity, including metabolites such as creatinine, valine, phospholipids in very large high-density lipoproteins HDL, triglycerides in small low-density lipoproteins (LDL), and apolipoprotein B, as well as lipids such as specific sphingomyelins (SM), phosphatidylcholines (PC), triacylglycerols (TAG) and diacylglycerols (DAG). By including any number of omics datasets and taking into account non-linear correlations, our method offers a truly system-level overview of biological pathways contributing to a comorbidity.
Methods
The Cardiovascular Risk in Young Finns Study
Our proposed method was applied to the 2007 wave of the Young Finns Study (YFS) dataset. This cross-sectional follow-up study presented a multidimensional array of data that encompasses (1) CVD-related phenotypes, depressive symptoms, and related risk factors, (2) metabolomic data, and (3) lipidomic data. This multidimensional dataset enables the identification of biomarkers, and the assessment of the effects of risk factors on both CVD and depression. Notably, it contains two omics datasets, allowing for multipartite projection to calculate projected scores between diseases.
For each participant, phenotype variables consist of 17 CVD-related indicators and six risk factors (see Table 1 and  4). Depressive variables include Beck’s score, and 21 individual depressive symptoms that contribute to the calculation of the summary score (see Table 2). The two omics datasets comprise 228 metabolomic variables and 437 lipidomic variables saved in two separate tables.
Missing data imputation and data discretization
We used random sample imputation to replace each missing value of a variable with a random sample drawn from existing values of that variable. This is conservative in the sense that it cannot introduce spurious correlations, and the more missing values a variable has, the lower its correlations. Following that, we retained the data points in which data of all omics variables were available, resulting in a total of 1686 participants (983 women, 703 men). The continuous variables were discretized by binning the data using a quantile-based method, dividing each continuous variable into probabilistically equal-sized buckets. Discretization is needed because part of the data is discrete and information-theoretic metrics cannot be compared for mixed data. The optimal number of quantiles was determined by Sturges’ rule. Categorical and discrete variables are included unchanged.
Significant MI correlation network
Mutual information (MI) was calculated on the discretized data to quantify the correlation between the YFS variables. This information-theoretic approach allows for capturing both linear and nonlinear relationships. It measures the average amount of information communicated in one random variable $X$ about another $Y$  [18].
$MI(X;Y)=\sum_{y\in Y}\sum_{x\in X}p(x,y)\cdot log\frac{p(x,y)}{p(x)p(y)},$ (1)
where $x$ and $y$ are the observations of variables $X$ and $Y$, $p(x)$ and $p(y)$ are marginal probabilities, $p(x,y)$ is the joint probability mass function.
Before constructing the significant MI network, we first applied a filtering process to minimize double counting and reduce redundant information among variables. The step was necessary because our proposed method incorporates the strength of MI correlations into the projection, and ensuring that each correlation is uniquely represented without redundancy is crucial for the method’s effectiveness. Consequently, some repeated variables and one variable from each highly correlated pair based on MI correlation were excluded from the analysis (see S1.1 in the Supplementary Information). This included Beck’s score, CVD-related variables (such as "fmd40", "fmd40pr", "fmd60", "fmd60pr", "fmd80", "fmd80pr", "maxfmd", "imtmax", "bbmax" and "volscore"), and another 115 variables in metabolites and lipids. This process resulted in a final dataset of 584 variables for further analysis.
We employed the bootstrapping method to compute $p$-values for MI correlations. Only correlations with $p$-values smaller than the significance level ($\alpha=0.01$) were included in the network, resulting in a significant multipartite MI correlation network. The corresponding adjacency matrix of the significant correlation network is denoted as $F(X;Y)$, in which each element is $f(X;Y)$.
The MI can quantify correlations directly between variables of phenotypes, symptoms and risk factors. However, these direct correlations would not be able to suggest any plausible causative mechanisms (pathways) between them, let alone identify key biomarkers. Our multipartite projection method computes these correlations indirectly as projected scores, based on their MI correlations with intermediate biomarkers.
Multipartite projection method
In the multipartite projection method, we defined the projected score as the sum of the average MI correlations between each pair of variables and their shared neighboring nodes (intermediate biomarkers). The definition is formulated as follows:
$w(X_{i};X_{j})=\sum_{k}\frac{f(X_{i};Y_{k})+f(X_{j};Y_{k})}{2},$ (2)
Where $X_{i}$ and $X_{j}$ are risk factor and/or phenotype variables and $Y_{k}$ is an intermediate biomarker that is a shared neighboring node of $X_{i}$ and $X_{j}$ in the tripartite MI network. Both $f(X_{i};Y_{k})$ and $f(X_{j};Y_{k})$ are non-zero.
This projection definition considers intermediate biomarkers in a weighted manner. When projected to different types of biomarkers this leads to a weighted, multilayer disease network of CVD phenotypes, depressive symptoms and related risk factors, in which each layer of the network corresponds to the type of biomarker involved in the projection. A stylized representation of this projection definition using average MI correlation is depicted in Figure 1. We also explored alternative definitions of the projected score, and selected this particular definition based on a comparative analysis (see Section S2.1 in the Supplementary Information).
Contribution of biomarkers to projected scores
The projection method allows us to calculate the contribution of each biomarker (in our case, metabolites and lipids) to each projected score. Considering the projection definition used in the main analysis and our focus on CVD-depression comorbidity, we defined the total contribution score of a biomarker to all possible projected links between CVD phenotypes and depressive symptoms as follows:
$CON(Y_{k})=\sum_{i,j}\frac{f(X_{i};Y_{k})+f(X_{j};Y_{k})}{2}$ (3)
where $Y_{k}$ is a biomarker, $X_{i}$ and $X_{j}$ represent pairs of phenotypes and symptoms. Both $f(X_{i};Y_{k})$ and $f(X_{j};Y_{k})$ are non-zero. Such a contribution score enables the identification of significant biomarkers that contribute to the overall comorbidity between CVD and depression. For a given pair of $(X_{i},X_{j})$, the method identifies the significant biomarkers that contribute to the specific projected score.
The relative importance of risk factors related to CVD and depression
It is widely recognized that CVD and depression are often linked to risk factors [19, 12]. When risk factors are included in the projection, the projected network provides a means of assessing the relative importance of risk factors in relation to a specific phenotype. The relative importance of a risk factor with respect to a particular phenotype is defined as the ratio of the projected score between the phenotype and the risk factor to the total projected score between the phenotype and all risk factors. This can be formulated as:
$r(Z_{i},X_{j})=\frac{w(Z_{i};X_{j})}{\sum_{i}w(Z_{i};X_{j})}$ (4)
where $X_{j}$ is a phenotype and $Z_{i}$ is one of the risk factors that has a link to the phenotype. $w(Z_{i};X_{j})$ is the projected score between the phenotype and the risk factor. The relative importance of risk factors considering CVD and depression is computed as the average of $r(Z_{i},X_{j})$ over all CVD-related phenotype and depressive symptoms, respectively. For a particular strong risk-disease link, we can identify the significant biomarkers associated with the relationship using Equation  3.
Results
Significant tripartite MI correlation network and projected multilayer networks
The projection method builds on a significant multipartite MI correlation network. Figure 2A displays the multipartite (tripartite, in our case) network, in which variables are partitioned into three groups: metabolome (blue), lipidome (green), and the combined group of phenotypes/symptoms of CVD (yellow) and depression (red) along with risk factors (purple). This network is tripartite because it shows only the links between metabolites/lipids and variables in the combined group. The edge width is determined by the mean MI value, calculated over 20 runs of random sample imputation. The node size is proportional to its weighted degree. To enhance readability and visualization, nodes in the tripartite network were further filtered based on their weighted degree, retaining 105 nodes and 669 edges. No depressive symptom variables appear in this network, as these variables exhibit limited correlations with biomarker variables. It is important to note that this filtering process is only for the purpose of network visualization; the projection analysis was performed on the network before filtering. Therefore, the depressive symptoms were still included in the analysis.
Observing the tripartite network, risk factors such as sex and BMI, CVD-related phenotypes like the ideal CVH score, average diastolic, and systolic blood pressure, exhibit a high weighted degree. This implies their extensive correlations with biomarker variables. However, based solely on this network, it is difficult to understand which biomarkers are driving these strong correlations, thereby obscuring the underlying causative mechanisms and biological pathways.
Figure 2B and C illustrate the metabolomic layer and lipidomic layer of the projected network of CVD-related phenotypes and depressive symptoms, respectively. In these projected networks, the link (edge) width is quantified by the mean projected score over 20 runs of random sample imputation. The node size is proportional to its weighted degree. We only include the projected links between CVD and depression variables due to our focus on the comorbidity between these two conditions. The projected networks are thus displayed in a bipartite layout. In the metabolomic layer of the projected network, we find strong projected links between five individual depressive symptoms (pessimism (’b2’), guilty feelings (’b5’), crying (’b10’), changes in appetite (’b18’), loss of libido (’b21’)) and three CVD-related variables (average diastolic blood pressure (’dkv’), average systolic blood pressure (’syst’), and ideal CVD health score (’idealCVH’)). In the lipidomic layer, strong projected links are observed between three depressive symptoms (’b10’, ’b18’ and ’b21’) and the same three CVD-related phenotypes. These strong projected links suggest a high risk of co-occurrence between these CVD-related phenotypes and depressive symptoms, predicting multimorbid phenotypes of CVD and depression.
Figure 3 compares the values of the MI correlation between CVD and depression with their projected score values from the projection method, showing a roughly positive linear relationship on a log-log scale. This pattern suggests that the projection method is effective, yet there remains a substantial degree of variability. We anticipate that this variability arises from the fact that the analysis includes only two omics of biomarkers. We expect that including additional categories of biomarkers would improve the correlation further.
Mediating biomarkers related to the comorbidity between CVD and depression
To investigate the CVD-depression comorbidity, we calculated the mean total contribution score of each metabolite and lipid to all projected links between CVD-related phenotypes and depressive symptoms over 20 runs of random imputation. Figure 4A and B display the top ten mediating metabolites and lipids that contribute to these projected CVD-depression links. The top mediating metabolites include creatinine, valine, leucine, phospholipids in very large HDL and small VLDL, triglycerides in small LDL and very small VLDL, and apolipoprotein B. The top mediating lipids consist of several specific sphingomyelins (SM), phosphatidylcholines (PC), triacylglycerols (TAG), diacylglycerols (DAG), phosphatidylglycerols (PG), and phosphatidylethanolamines (PE). These metabolites and lipids were identified as significant biomarkers that link CVD and depression through biological pathways such as inflammation, cholesterol regulation, and atherosclerosis. Most of these metabolites and some lipids are already known, while others, such as SM and PG, are potentially novel biomarkers. Detailed rankings and further information on these biomarkers can be found in Table 3.
Relative importance of risk factors and related biomarkers
Including risk factors in the projection results in a projected multilayer network, in which nodes are partitioned into three groups: risk factors, CVD-related phenotypes, and depressive symptoms (see Figure 5A and B). Observing both metabolomic and lipidomic layers of the projected network, the high weighted degree of sex and BMI indicates their extensive associations with both CVD and depression compared to other risk factors. The relative importance of all risk factors is shown in Table 4. The results suggest that sex and BMI are two important risk factors related to both CVD and depression, which corroborates previous studies [20, 21, 22, 23]. It is important to note that sex can be a covariate or confounder. Therefore, we conducted gender-specific analyses (see S2.3 in the Supplementary Information). Additionally, age–another well-known risk factor–does not show high importance in the YFS dataset, likely because the YFS participants were all young adults.
We took two demonstrated high-score projected links ("Sex"–"Appetite (’b18’)" and "Sex"–"idealCVH") as examples to further analyze the contributing biomarkers. In this case, sex may act as a confounding factor, influencing both variables. Therefore, it is plausible for appetite and the ideal CVH score to exhibit a strong projected score in the projected multilayer networks (see Figure 2). The significant mediating biomarkers based on the mean total contribution score are shown in Figure 5C, D. Details of these biomarkers can be found in Table S6 and S7. For the link between sex and appetite, significant metabolites include creatinine, free cholesterol in large HDL, and phospholipids, cholesterol and triglycerides in very large VLDL (very low-density lipoprotein). Significant lipids are mainly sphingomyelins (SM) 32:2 and lysophosphatidylcholines (LPC). For the link between sex and the ideal CVH score, significant metabolites encompass phospholipids, free and total cholesterol in large HDL and LDL, apolipoprotein B, and Leucine. Significant lipids are primarily triacylglycerols (TAG) and diacylglycerols (DAG). These identified biomarkers shed light on how risk factors influence the development of CVD and depression.
Discussion
In this study, we proposed a multipartite projection method to infer the relationships between CVD and depression, leading to the creation of a weighted multilayer network of CVD-related phenotypes and depressive symptoms. The proposed method can be considered a generalization of previous disease network modeling approaches [13, 15, 16]. It demonstrates several advantages: (i) it uses MI correlation to capture both linear and nonlinear associations; (ii) it is capable of integrating multiple omics datasets into the projection; (iii) it defines a refined projection strategy that accounts for shared biomarkers considering their correlations in a weighted way. Additionally, this method has the potential to incorporate directed causal relationships and higher-order interactions (e.g. synergistic relationship) into the projection. It also has the capability to construct weighted higher-order projected disease networks by considering the shared neighboring nodes of more than two disease variables.
We applied the method to the YFS dataset and inferred the projected scores between CVD-related phenotypes and depressive symptoms considering two available omics datasets: metabolome and lipidome. A weighted multilayer disease network was created based on the projected scores. Each layer of the projected network represents the links between diseases in a specific omics context, which is metabolomic or lipidomic in this study. This framework allows for a rich network representation of CVD and depression across multiple dimensions and facilitates the identification of related biomarkers. It provides better insights into possible causal mechanisms and biological pathways underlying the links between these diseases, getting closer to explaining comorbidities.
The most contributing metabolites and lipids (sorted in Figure 4) are inferred to play a role in the biological pathways connecting CVD and depression. In metabolites, Creatinine is a byproduct of breaking down creatine. On the depression side, creatine presents an antidepressant-like effect [24, 25]. On the CVD side, creatinine is associated with several traditional cardiovascular risk factors such as diabetes and dyslipidemia in hypertensive patients and stroke [26]. Valine has been considered to be significantly associated with depression in recent studies [27, 28]. High levels of branched chain amino acids (BCAAs), which both valine and leucine are categorized as, predicted a higher risk of developing CVD [27, 29]. Such particles are also rich in phospholipids which are known to play a crucial role in protecting against cardiovascular disease by promoting cholesterol efflux and reducing inflammation [30, 31]. The anti-inflammatory effects of phospholipids within HDL may simultaneously improve depressed mood and increase treatment response to conventional antidepressant medications in patients with major depression [32]. Such an anti-inflammatory property is also associated with apolipoprotein A-I, a major protein component of HDL particles [33]. This may well explain why "ApoBApoA1" is also in the ranking. Further, high cholesterol has been recognized as one of the significant risk factors for cardiovascular disease (CVD) [34]. It also shows a positive correlation with depression severity [35].
On the lipid side, the projection method identified lipids such as several specific SM, PC, TAG, DAG, PG and PE as significant biomarkers (Figure 4). PC and PE are two major classes of phospholipids that have been associated with CVD and depression. They play important roles in numerous biological processes and are the most abundant phospholipids in mammalian cell membranes. Their levels have been linked to CVD risk and outcomes [36]. Alterations in PC and PE, have been observed in individuals with psychotic experiences, which are associated with an increased risk of mental disorders, including depression [37]. Additionally, TAG and DAG have also been reported to be associated with both CVD and depression in many studies [38, 39, 40, 41].
Notably, the projection method identified SM 32:2, SM 40:0 and PG 36:1, as significant biomarkers associated with both CVD and depression. While direct evidence linking these biomarkers to CVD and depression is limited, exploring novel biomarkers like SM and PG in the research of CVD-depressioin comorbidity could offer new insights into the underlying biological pathways. Currently, SM 32:2 (SM(d18:2/14:0)) shows a significant and positive association with atherosclerosis, a key risk factor for CVD [42]. Additionally, some SM species have been reported to be linked with depression and anxiety [43, 44]. PG is an intermediate in the biosynthesis of various lipids, particularly cardiolipin (CL) [45, 46], which may be associated with mitochondrial function and cardiac health [47]. However, further research is needed to elucidate the precise role of SM and PG in the comorbidity between CVD and depression.
When turning our attention to risk factors, the projected network identified sex and BMI as two important risk factors linking CVD and depression. Sex is indeed a risk factor (e.g., covariate, confounder) for these health conditions [48, 12]. Men, on average, are more prone to certain risks, such as high blood pressure, higher levels of LDL (bad) cholesterol, and a higher likelihood of smoking and excessive alcohol consumption. Consequently, the risk of developing CVD tend to begin earlier in men, often around middle age. Conversely, women are more likely to be diagnosed with depression [49]. The sex-specific difference in depression can be attributed to the sex-induced biological changes such as in hormones in women, and different presentations of depressive symptoms [50]. It is important to note that sex can act as a covariate or confounder. Thus, we conducted gender-specific analyses using the projection method. Further discussion on these results is provided in S2.3 in the Supplementary Information.
BMI is considered to be highly associated with both CVD and depression. High BMI, particularly when it falls into the obese category (BMI of 30 or higher), is a well-established risk factor for cardiovascular diseases [51, 52]. Depression shows a significant U-shaped association with BMI categories (underweight, normal, overweight and obesity) [53, 54]. In addition, age is another important risk factor associated with both CVD and depression [55], but not in the YFS as the participants in this study were young adults.
The biomarkers underlying the strong links between risk factors and phenotypes were detected. These biomarkers are considered important in elucidating how risk factors contribute to disease development and understanding the underlying biological pathways involved. For instance, creatinine is a gender-specific metabolite [56, 57], and has been recognized as a nutritional marker associated with decreased appetite [58, 59]. SM 32.2 and many specific LPCs are identified as biomarkers related to appetite in individuals with depression, although no human studies have reported this finding. However, a recent paper using mouse models discovered that an increase in LPCs in the blood following a prolonged fasting increased lysophosphatidic acid (LPA) level in the brain [60]. These LPAs can modulate the excitability of nerve cells in the cerebral cortex, thereby influencing food intake and appetite regulation. Additionally, the detection of TAG and DAG in the context of sex-related cardiovascular research suggests potential lipid-related mechanisms underlying sex disparities in CVD susceptibility and outcomes [61]. TAG concentrations are consistently lower in healthy lean and obese women than in men [62], and its elevated level has been reported to be highly associated with CVD [63, 64, 38].
This study has several limitations. Firstly, we conducted two steps to eliminate redundant information based on MI correlation. However, this process might lead to an underestimation of projected scores, as some intermediate biomarkers that are not entirely redundant may have been omitted. Secondly, the YFS dataset includes only limited omics data: metabolome and lipidome, which may not be sufficient for a complete construction of the projected links. These two limitations likely contribute to the consistently low explaining score (see S2.1 in the Supplementary Information) for all projection definitions. This finding underscores the importance of including multiple omics data in the projection. Another limitation pertains to the absence of directionality in the MI correlation. MI is a correlation measure rather than a causal discovery method, which makes it incapable of capturing causal relationships between variables. It is therefore challenging to determine whether a biomarker or risk factor plays a role of a confounder, collider, mediator or moderator. Finally, our method is blind to so-called higher-order correlations or synergistic associations beyond pairwise relationships. This is a common limitation in the field, as similar approaches in the literature have the same constraint.
This study may provide direction for future data gathering efforts towards biomarkers that are predicted as highly important unexpectedly such as specific SM and PG. Future work should also involve incorporating directed causal [65] and higher-order relationships [66] into the projection and replicating the analysis across diverse datasets, including more omics data such as genomics and proteomics.
Conclusions
In this study, we introduced a novel multipartite projection method to construct multilayer disease networks, focusing on the comorbidity between CVD and depression. This approach complements traditional biological and experimental analyses by offering insights into the underlying biological pathways. By leveraging the MI correlation, our method integrates multiple omics datasets, such as metabolomics and lipidomics, to provide a system-level overview of the biological pathways contributing to this comorbidity.
Our analysis, applied to data from the Young Finns Study, identified significant biomarkers that mediate the relationship between CVD and depression. These biomarkers include metabolites like creatinine, valine, and phospholipids in very large high-density lipoproteins (HDL), as well as lipids such as sphingomyelins (SM), phosphatidylcholines (PC), Phosphatidylglycerols (PG), triacylglycerols (TAG), and diacylglycerols (DAG). Additionally, we highlighted sex and body mass index (BMI) as important risk factors influencing the comorbidity.
The proposed method addresses several limitations of previous approaches by incorporating multiple omics layers, accounting for non-linear correlations, and providing a weighted measure of biomarker contributions. This allows for a more nuanced understanding of disease associations and potential pathways.
The findings of this study underscore the importance of a multidimensional approach in studying complex diseases. The multipartite projection method offers a powerful tool for uncovering the intricate biological mechanisms underlying comorbid conditions, paving the way for more targeted preventive and therapeutic strategies. Future research could expand this methodology to other disease pairs and incorporate additional omics datasets to further enhance our understanding of disease comorbidities.
Funding
The author(s) declare financial support was received for the research, authorship, and/or publication of this article. Jie Li, Jos A. Bosch and Rick Quax were supported by the EU project To_Aition, which has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 848146. Cillian Hourican and Rick Quax were supported by the Netherlands Organisation for Health Research and Development (ZonMw), Open Competition Grant 09120012010063. The Young Finns Study has been financially supported by the Academy of Finland: grants 356405, 322098, 286284, 134309 (Eye), 126925, 121584, 124282, 129378 (Salve), 117797 (Gendi), and 141071 (Skidi); the Social Insurance Institution of Finland; Competitive State Research Financing of the Expert Responsibility area of Kuopio, Tampere and Turku University Hospitals (grant X51001); Juho Vainio Foundation; Paavo Nurmi Foundation; Finnish Foundation for Cardiovascular Research; Finnish Cultural Foundation; The Sigrid Juselius Foundation; Tampere Tuberculosis Foundation; Emil Aaltonen Foundation; Yrjö Jahnsson Foundation; Signe and Ane Gyllenberg Foundation; Diabetes Research Foundation of Finnish Diabetes Association; EU Horizon 2020 (grant 755320 for TAXINOMISIS and grant 848146 for To Aition); European Research Council (grant 742927 for MULTIEPIGEN project); Tampere University Hospital Supporting Foundation; Finnish Society of Clinical Chemistry; the Cancer Foundation Finland; pBETTER4U_EU (Preventing obesity through Biologically and bEhaviorally Tailored inTERventions for you; project number: 101080117); and the Jane and Aatos Erkko Foundation. PPM was supported by the Academy of Finland (Grant number: 349708); M.A.-K. was supported by a research grant from the Sigrid Juselius Foundation, the Finnish Foundation for Cardiovascular Research, and the Research Council of Finland (grant no. 357183).